HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

QUANTUM TUNNELLING IN MOLECULES AND CLUSTERS

\underline{J. O. RICHARDSON} , Department of Chemistry and Applied Biosciences, ETH Zurich, Switzerland

Quantum tunnelling leaves fingerprints in the high-resolution spectra of molecules and clusters. This provides a serious challenge for theory to predict and interpret the spectrum. I will discuss recent progress in methods based on path-integral formalism of quantum mechanics, which leads to practical methods for simulating tunnelling even in reasonably large molecules. Instanton theory is based on a semiclassical approximation, from which one obtains a powerful approach requiring no more than a single classical trajectory.^{1} The method has been used to elucidate the tunnelling splitting patterns in many molecules including water clusters.^{2} More recently, we have developed perturbative corrections to instanton theory which can account for anharmonicity and correct the results in the case of low barriers.^{3} For extremely anharmonic systems such as methyl rotations, a nonperturbative solution is required. We have therefore developed a path-integral molecular dynamics approach, which in principle gives identical results to an exact solution of the Schrödinger equation and employs an “Eckart spring” to rigorously project out the ground rotational state. However, our approach does not require wavefunctions and thus avoids the curse of dimensionality of traditional quantum-mechanical approaches, but instead scales linearly with system size.^{4}

110.1063/1.5028352, J. O. Richardson, Perspective: Ring-polymer instanton theory, J. Chem. Phys.

148, 200901 (2018).

210.1126/science.aae0012, J. O. Richardson, C. Pérez, S. Lobsiger, A. A. Reid, B. Temelso, G. C.

Shields, Z. Kisiel, D. J. Wales, B. H. Pate and S. C. Althorpe, Concerted Hydrogen-Bond Breaking by Quantum Tunneling in the Water Hexamer Prism, Science 351, 1310-1313 (2016).

3arXiv:2304.10963, J. E. Lawrence, J. Dušek and J. O. Richardson, Perturbatively corrected ringpolymer instanton theory for accurate tunneling splittings, accepted J. Chem. Phys. (2023).

4arXiv:2305.11002, G. Trenins, L. Meuser, H. Bertschi, O. Vavourakis, R. Flütsch and J. O. Richardson, Exact tunnelling splittings from symmetrized path integrals, accepted J. Chem. Phys. (2023).

p-number: p247  Submitted on Mon, 26 Jun 2023 10:25:44 +0200